OpenStudy (anonymous):

Solve the PDE: df/dx+2x*df/dy=y d=\partial, lazy. :)

OpenStudy (anonymous):

Suggested substitution is u=ax^2+y,v=x.

OpenStudy (anonymous):

I arrive at \[\frac{\partial f}{\partial v}=y\]not sure where to go from here though.